# What is ‘Typological Diversity’ in NLP?

### v1

@@ Abstract
The NLP research community has devoted increased attention to languages beyond English, resulting in considerable improvements for multilingual NLP. However, these improvements only apply to a small subset of the world’s languages. Aiming to extend this, an increasing number of papers aspires to enhance generalizable multilingual performance across languages. To this end, linguistic typology is commonly used to motivate language selection, on the basis that a broad typological sample ought to imply generalization across a broad range of languages. These selections are often described as being ‘typologically diverse’. In this work, we systematically investigate NLP research that includes claims regarding ‘typological diversity’. We find there are no set definitions or criteria for such claims. We introduce metrics to approximate the diversity of language selection along several axes and find that the results vary considerably across papers. Furthermore, we show that skewed language selection can lead to overestimated multilingual performance. We recommend future work to include an operationalization of ‘typological diversity’ that empirically justifies the diversity of language samples.
[…]
github.com/WPoelman/typ-div
@@
What is ‘Typological Diversity’ in NLP?
@@ 1 Introduction
Figure 1: Number of papers with ‘typological diversity’ claims published by year.
@@
This data sparsity can be mitigated by leveraging cross-lingual transfer through the training of a language model on multilingual data. Despite the potential of multilingual language modelling, common methodologies are primarily developed for English. But there is no guarantee that an approach that works well for one language will work equally well for others (Gerz et al., 2018). For instance, morphologically complex languages can be over-segmented by current widely-used tokenization methods (Rust et al., 2021). Evaluation on a broad range of languages is important for drawing more generalizable conclusions about the performance of multilingual language technology. For instance, including only morphologically simple languages, such as English, can give an unrealistic image of the effectiveness of a tokenization method, because morphologically simple languages are generally easier to tokenize compared to complex ones.
@@
Current work increasingly evaluates models on multiple languages, but because of resource constraints, it is not realistic to test a model on the thousands of languages in the world. In order to still ensure a degree of generalizability, previous work recognizes the importance of diverse language sampling. Ponti et al. (2020) suggest that merely evaluating on a small set of similar languages is an unreliable method for estimating a multilingual model’s performance, since such evaluations lack robustness to previously unseen typological properties, for instance. For mitigating these sampling issues, NLP research increasingly relies on linguistic typology (see Figure 1). Evaluating on a set of so-called ‘typologically diverse’ languages implies robustness over typological variations, implicitly claiming generalizability across languages more generally. Importantly however, there is a fundamental lack of common understanding of exactly what claims of typological diversity entail.
@@
In this work, we systematically investigate how the term ‘typological diversity’ is used in NLP research, by surveying the literature for past and current practices. We find that there is no set definition of ‘typological diversity’ in NLP: justifications for claims range from (i) none whatsoever, to (ii) using language phylogeny as a proxy for typology, to (iii) basing diversity on a more or less well-founded subset of available typological features. As a consequence, typological language similarity and typological feature value inclusion vary considerably across papers (Section 5). Furthermore, we demonstrate why this is problematic, by showing how a skewed selection of evaluation languages can lead to overstated multilingual performance and poor generalizability (Section 6). We recommend future approaches that use the term ‘typologically diverse’ to document their language selection, and to add a measure of typological diversity, such as average language distance (Section 5.1) or typological feature inclusion (Section 5.2). Better-documented language sampling can be crucial for the development of multilingual NLP, because it provides an insight into commonly overlooked linguistic difficulties of multilingual language modelling.
2 Related work
@@
Language sampling techniques are used in linguistic typology as well as in NLP. Although these fields both concern languages and generalizability, they typically have different objectives and means, which we contrast in this section.
@@
Language sampling has a long-standing history in linguistic typology. Rijkhoff et al. (1993) state that testing a linguistic hypothesis on a few similar languages is not sufficient for drawing conclusions on human language as a whole. Rather, they argue that generalizable samples should capture the diversity of the world’s languages. Language samples can be constructed in a number of ways. Rijkhoff and Bakker (1998) distinguish three types of sampling: random sampling, probability sampling and variety sampling. Random samples do not have any specific stratification criteria, whereas probability samples are constructed to minimize bias, for instance by avoiding a skew towards languages from particular language families or geographical areas. Lastly, variety samples are designed to contain as much linguistic variety as possible, including relatively uncommon typological characteristics. The following quote from Miestamo et al. (2016) on variety sampling provides a useful summary of this language sampling perspective:
The goal of variety sampling is to maximize the amount of variation in the language data. Obviously, [this] can be covered by including every one of the world’s approximately 7,000 languages (…) However, this is not possible for a number of reasons: not nearly all languages have been sufficiently described (…), even if sources exist, they might not be accessible to the researcher; and even if one had access to adequate sources for all 7,000 languages, it would (in the case of most research questions) be too time-consuming to include them all. Variety sampling is therefore needed.
@@
Language sampling methods in NLP have a similar goal. That is, to sample languages in such a way that the findings (e.g., multilingual performance estimates) can reasonably be expected to generalize across languages. Yet, their justifications are commonly less rigorous than those in linguistic typology. For language sampling methods in NLP, it is usually not explicitly mentioned which type of sampling is conducted, and for which reasons. In practice, most papers do not provide a justification for their selections (see Section 4.3), which resembles random sampling. A rare exception to this is Ponti et al. (2020), who use variety sampling and mention that this ensures evaluation that is robust towards typological characteristics that were probably not seen during training.
Means
@@
While the goal of language selection for ensuring generalizable conclusions is similar between NLP and linguistic typology, the ideal means are typically different. In linguistic typology, the variables under investigation are often typological characteristics. A typologist may aim to test whether a correlation of features holds across languages. Such a correlation could then be referred to as a ‘universal’: languages with object-verb ordering would be postpositional Greenberg (1963), for instance. To avoid circularity, language sampling is not done directly based on those features. Instead, popular methods rely on (combinations of) language family, genus, and geography (Miestamo et al., 2016).
@@
By contrast, what is evaluated in multilingual NLP does not concern typological features directly, but rather, model performance across languages. In fact, to support generalizability claims, multilingual benchmarks must explicitly contain a range of typological phenomena. Relying on geographical and genetic relations between languages is not guaranteed to ensure this. Nichols (1996) argues that “the kind of variables that define genealogical groups and tree shapes have a very different nature from the kind of variables that define typological diversity.” For estimating the generalizability of a language sample in NLP evaluation, it is possible, and even desired, to have direct insight into the actual typological properties that are included in the selection (see also Section 6).
@@ 3.1 Data Collection
We retrieve papers from the ACL Anthology (*ACL, COLING, LREC, etc.) and from several top-tier AI venues where NLP research is published (NeurIPS, AAAI, ICLR, IJCAI and ICML). [^1] See Appendix A for details. From these venues, we select papers that contain the following search string in either their title or abstract:
@@
Our search string aims to capture a broad range of formulations, including: “typologically and genetically diverse languages” Xu et al. (2020), “languages of diverse typologies” Eskander et al. (2022) and “diverse languages in terms of language family and morphological typology” Eskander et al. (2020b). Another commonly used phrase is “we evaluate on $N$ typologically different languages”. While we acknowledge that some researchers might (unintentionally) use the two claims interchangeably, we only focus on claims regarding ‘diversity’. It is reasonable to indicate that two languages are ‘typologically different’, even though they might be very similar (e.g., Dutch and German).
@@ 3.2 Annotation Guidelines
We annotate the retrieved papers according to the following criteria:
@@
For criterion 1, we only consider claims made by the paper at hand. That is, if a paper uses a dataset which is claimed to be ‘typologically diverse’, but the current paper does not mention this, we do not consider this to be a claim. This also applies to related work: if a paper references a claim made by another paper, but they themselves do not, then we do not consider this as a claim. The reason we annotate whether a paper introduces a new dataset, is that datasets are often starting points for subsequent projects. If datasets are claimed to be ‘typologically diverse’, this claim is spread not only through the paper introducing it, but also the dataset itself. An overview of all papers introducing a dataset for ‘typologically diverse’ languages can be found in Appendix C, Table 3.
Lastly, we annotate which languages are used in the papers making a claim. These are normalized to ISO-639-3 codes if available. When applicable, we only select the languages about which the claim is made. For example, Kann et al. (2020) and Shi et al. (2023) specifically distinguish between baseline languages and their typologically diverse test language selection.
@@ 3.3 Inter-Annotator Agreement
In order to determine the agreement between the annotators, and estimate the quality of the annotation guidelines, we calculate Cohen’s $\kappa$, as there are two annotators and we treat the values as binary. The agreement values of all annotation items are shown in Table 1. For the annotations concerning claims (1) and datasets (2), we achieve substantial agreement. We resolve the disagreements by discussing the annotations and merge them into a single label. For the sake of transparency, and to support incentives for human label variation Plank (2022), we also release the non-aggregated annotations. The agreement for the language selection annotation (3) is somewhat low, which is partially explained by the fact that we calculate agreement over all language together, rather than on the micro-level for individual languages. This is because there is no ground truth regarding the number of languages to annotate. Furthermore, a number of inconsistencies there are due to ISO-639-3 variants (jap $\rightarrow$ jpn or ger $\rightarrow$ deu) and different codes for ambiguous language mentions (just ‘Norwegian’ could be nor or nob). We iteratively resolve these issues to get our final list of languages per paper.
@@ 4 Results
4.1 Publication Overview
In total we retrieve 194 papers, of which 110 contain a claim. The rest of this analysis is based on these 110 papers. Figure 2 shows the most common venues where papers with a claim have been published. The top six venues are all high-ranking NLP conferences, with SIGMORPHON being the first workshop at the seventh spot. The AI-centered venues only have a small number of papers with claims; ICLR and AAAI both have two, ICML one, and the rest none. In total, 38 papers introduce a dataset, [^2] Details are listed in Appendix C, Table 4. most of which are published at ACL and LREC since 2020 (Figure 1 and 2).
Figure 3: Number of papers using $N$ languages.
@@ 4.2 Language Overview
Four of the papers that claim typological diversity do not mention which languages they use at all. The other papers use a median of 11 languages, with a minimum of 2 and a maximum of 90. In total, the papers use 315 unique languages, of which 160 are used just once. Figure 3 shows the distribution of the number of languages used per paper.
English is the most used language (63 papers), followed by German (60), Russian (58) and Finnish (57). [^3] The long tail of this distribution is listed in Appendix B, Figure 8. Most papers use between 5 and 20 languages. The paper that uses the most languages is Vylomova et al. (2020), which contains 90. Figure 4 shows where these languages are primary spoken according to the World Atlas of Language Structures (WALS) Dryer and Haspelmath (2013). We observe a skew towards languages spoken in Europe, where multiple languages are used in more than 50 papers, while not a single language from the Americas is included in more than 10 papers.
@@ 4.3 Justifications
Still, most other papers justify their typological diversity in a ‘post-hoc’ way. They do not mention the language selection considerations based on which languages were selected. Rather, they mention how diverse the sample is after it has been created. The XTREME-R dataset Ruder et al. (2021), which we will discuss further in Section 6, is exemplary of this. They mention ‘diversity indices’ that cover family, as well as typology and compare theirs to other datasets, but do not detail how or if these were used in sampling.
@@ 5 Typological Analysis
From Section 4.2, it follows that – when looking at all papers that claim ‘typologically diverse’ language selection – there is a geographical skew in language selection, where languages from certain areas are over-represented (Figure 4). A skew is not surprising, as it is often more feasible to select languages for which there are existing resources and typological descriptions, than languages for which these have to be gathered from scratch. This constitutes bibliographic bias (Rijkhoff and Bakker, 1998). However, we cannot draw conclusions about typological diversity on this basis, as typology is not equal to geography (Cysouw, 2013). Therefore, in this section, we analyze the typological diversity according to metrics that take into account syntactic features, and the absolute included typological feature values. We do this by comparing pairwise language distances (Section 5.1) and absolute typological feature value inclusion (Section 5.2) across papers. Lastly, we look into the relationship between the number of included languages and typological diversity, to help guide future language selection efforts (Section 5.3).
@@ 5.1 Mean Language Distance
We first approximate the typological diversity of each paper’s language selection with lang2vec (Littell et al., 2017). This toolkit provides language distances that are calculated based on language vectors from URIEL, a resource that contains information from a range of typological databases such as WALS and PHOIBLE. We choose URIEL and lang2vec specifically, because its aggregated information enables analysis across databases, and because its vectorized format mitigates issues stemming from incomplete feature coverage in typological databases.
For each paper in our survey that claims a typologically diverse language set and that is covered in lang2vec (312 / 315 unique languages), we calculate the mean pairwise syntactic distance (MPSD) for the included languages. That is, we take the average of all pre-computed syntactic lang2vec distances for each pair of languages in the selection. We only take into account languages that have at least 5% coverage in the URIEL vectors. We do this to make sure the very low coverage languages do not distort the results. This affects 31 papers in our dataset, resulting in a total of 242 unique languages. Figure 5 shows the distribution of MPSDs across papers. We illustrate the range of syntactic distance with some examples from well-covered languages. The distance between Danish and Norwegian (same family, same genus) is 0.22, Danish-Spanish (same family, different genus) is 0.59 and Danish-Japanese (different family, different genus) is 0.69. Additionally, we compare our typology-based approximations with mean geographical and genetic pairwise distances. Interestingly, the genetic distance is much higher and less spread out than the syntactic distance. This emphasizes that genealogical sampling does not by default ensure typological diversity.
The typological distances vary considerably, with outliers on either side (Figure 5). In papers claiming typological diversity, the minimum MPSD is found in Goel et al. (2022), who use English, French and Spanish. The maximum MPSD is found for North Sámi, Galician, and Kazah (Vania et al., 2019).

### v3

# What is “Typological Diversity” in NLP?

@@ Abstract
The NLP research community has devoted increased attention to languages beyond English, resulting in considerable improvements for multilingual NLP. However, these improvements only apply to a small subset of the world’s languages. An increasing number of papers aspires to enhance generalizable multilingual performance across languages. To this end, linguistic typology is commonly used to motivate language selection, on the basis that a broad typological sample ought to imply generalization across a broad range of languages. These selections are often described as being ‘typologically diverse’. In this meta-analysis, we systematically investigate NLP research that includes claims regarding typological diversity. We find there are no set definitions or criteria for such claims. We introduce metrics to approximate the diversity of resulting language samples along several axes and find that the results vary considerably across papers. Crucially, we show that skewed language selection can lead to overestimated multilingual performance. We recommend future work to include an operationalization of typological diversity that empirically justifies the diversity of language samples. To help facilitate this, we release the code for our diversity measures. [^1] Our code and data are publicly available: https://github.com/WPoelman/typ-div
@@
What is “Typological Diversity” in NLP?
@@ 1 Introduction
Figure 1: There is an increase in the number of publications with ‘typological diversity’ claims across time.
@@
This data sparsity can be mitigated by leveraging cross-lingual transfer through the training of a language model on multilingual data. Despite the potential of multilingual language modelling, methodologies are primarily developed for English. However, there is no guarantee that an approach that works well for one language will work equally well for others (Gerz et al., 2018). For instance, morphologically complex (or rich) languages can be over-segmented by current widely-used tokenization methods (Rust et al., 2021). Evaluation on a broad range of languages is important for drawing more generalizable conclusions about the performance of multilingual language technology Bender (2009, 2011); Pikuliak and Simko (2022). Evaluating a tokenization method on onlymorphologically simple languages, such as English, can give an unrealistic image of the effectiveness of a tokenization method, because morphologically simple languages are generally easier to tokenize compared to complex ones.
@@
Current work increasingly evaluates models on multiple languages, but because of resource constraints, it is not realistic to test a model on the thousands of languages in the world. In order to still ensure a degree of generalizability, previous work recognizes the importance of diverse language sampling. Ponti et al. (2020) suggest that merely evaluating on a small set of similar languages is an unreliable method for estimating a multilingual model’s performance, since such evaluations lack robustness to previously unseen typological properties, for instance. In an attempt to address these sampling issues, NLP research increasingly claims to rely on linguistic typology (see Figure 1). Evaluating on a set of so-called ‘typologically diverse’ languages implies robustness over typological variations, implicitly claiming broad generalizability across languages. Importantly however, there is a fundamental lack of common understanding of exactly what a claim of typological diversity entails.
@@
In this work, we systematically investigate how the term ‘typological diversity’ is used in NLP research by surveying the literature for past and current practices. We find that there is no set definition of ‘typological diversity’ in NLP: justifications for claims range from (i) none whatsoever, to (ii) using language phylogeny as a proxy for typology, to (iii) basing diversity on a more or less well-founded subset of available typological features. As a consequence, typological language similarity and typological feature value inclusion vary considerably across papers (§5). Furthermore, we demonstrate why this is problematic, by showing how a skewed selection of evaluation languages can lead to overstated multilingual performance and poor generalizability (§6). We recommend future approaches that use the term ‘typologically diverse’ to document their language samples, and to add a measure of typological diversity, such as the mean pairwise language distance (§5.1) or typological feature inclusion (§5.2). Better-documented language sampling is crucial for the development of multilingual NLP, because it provides insights into commonly overlooked linguistic difficulties of multilingual language modelling. We provide software to verify and visualize this, linked below the abstract.
2 Background
@@
Language sampling techniques are used in linguistic typology as well as in NLP. Although these fields are both concerned with languages and generalizability, they typically have different objectives and resources, which we contrast in this section.
@@
Language sampling has a long-standing history in linguistic typology. Rijkhoff et al. (1993) state that testing a linguistic hypothesis on a few similar languages is not sufficient for drawing conclusions on human language as a whole. Rather, they argue that generalizable samples should capture the diversity of the world’s languages. Language samples can be constructed in a number of ways. Rijkhoff and Bakker (1998) distinguish three types of sampling: random, probability and variety. Random samples do not have any specific stratification criteria, whereas probability samples are constructed to minimize bias, for instance by avoiding a skew towards languages from particular language families or geographical areas. Lastly, variety samples are designed to contain as much linguistic variety as possible, including relatively uncommon typological characteristics (Miestamo et al., 2016).
@@
Language sampling methods in NLP have a similar goal: to sample languages in such a way that the findings (e.g., multilingual performance estimates) are expected to generalize across languages. Yet, their justifications are commonly less rigorous than those in linguistic typology. For language sampling methods in NLP, it is usually not explicitly mentioned which type of sampling is conducted, and for which reasons. In practice, most papers do not provide a justification for their selections (see §4.3), which resembles random sampling. A rare exception to this is Ponti et al. (2020), who use variety sampling and mention that this ensures evaluation that is robust towards typological characteristics that were likely not seen during training.
Resources
@@
While the goal of language sampling for ensuring generalizable conclusions is similar between NLP and linguistic typology, their resources are typically different. In linguistic typology, the variables under investigation are often typological characteristics. A typologist may aim to test whether a correlation of features holds across languages. Such a correlation could then be referred to as a ‘universal’: languages with object-verb ordering would be postpositional Greenberg (1963), for instance. To avoid circularity, language sampling is not done directly based on those features. Instead, popular methods rely on (combinations of) language family, genus, and geography (Miestamo et al., 2016).
@@
By contrast, what is evaluated in multilingual NLP does not concern typological features directly, but rather, model performance across languages. In fact, to support generalizability claims, multilingual benchmarks must explicitly contain a range of typological phenomena (Bender, 2009, 2011; Pikuliak and Simko, 2022). Relying on geographical and genetic relations between languages is not guaranteed to ensure this. Nichols (1996) argues that “the kind of variables that define genealogical groups and tree shapes have a very different nature from the kind of variables that define typological diversity.” For estimating the generalizability of a language sample in NLP evaluation, it is possible, and even desired, to have direct insight into the actual typological properties that are included in the selection (see also §6).
@@ 3.1 Data Collection
We retrieve papers from the entire ACL Anthology (*ACL, COLING, LREC, etc.) and from several top-tier AI venues where NLP research is published (NeurIPS, ICLR, AAAI, etc.). [^2] See Appendix A for details. We select papers that contain the following search string in either their title or abstract:
@@
Our search string aims to capture a broad range of formulations, including: “typologically and genetically diverse languages” Xu et al. (2020), “languages of diverse typologies” Eskander et al. (2022) and “diverse languages in terms of language family and morphological typology” Eskander et al. (2020b). Another commonly used phrase is “we evaluate on $N$ typologically different languages”. While we acknowledge that some researchers might (unintentionally) use the two claims interchangeably, we only focus on claims regarding ‘diversity’. It is reasonable to state that languages are ‘typologically different’, even though they might be similar (e.g., Dutch and German).
@@ 3.2 Annotation Guidelines
The first two authors annotated the retrieved papers according to the following criteria:
@@
For criterion 1, we only consider claims made by the paper at hand. That is, if a paper uses a dataset which is claimed to be ‘typologically diverse’, but the current paper does not mention this, we do not consider this to be a claim. This also applies to related work: if a paper references a claim made by another paper, but they themselves do not make a claim, then we do not consider this as a claim. The reason that we annotate whether a paper introduces a new dataset is that datasets are often starting points for subsequent projects. If datasets are claimed to be ‘typologically diverse’, this claim is spread not only through the paper introducing it, but also the dataset itself. [^3] An overview of all papers introducing a dataset for ‘typologically diverse’ languages is in Appendix D, Table 3.
Lastly, we annotate which languages are included in the papers making a claim. These are normalized to ISO-639-3 codes if available. Where applicable, we only select the languages about which the claim is made. For example, Kann et al. (2020) and Shi et al. (2023) specifically distinguish between baseline languages and their typologically diverse test language selection.
@@ 3.3 Inter-Annotator Agreement
In order to determine the agreement between the annotators, and estimate the quality of the annotation guidelines, we calculate Cohen’s $\kappa$, as there are two annotators and we treat the values as binary. The agreement values of all annotation items are shown in Table 1.
@@
For the annotations concerning claims (1) and datasets (2), we achieve substantial agreement. We resolve the disagreements by discussing the annotations and merge them into a single label. For the sake of transparency, and to support incentives for human label variation Plank (2022), we also release the non-aggregated annotations. [^4] At the repository listed below the abstract. The agreement for the language selection annotation (3) is somewhat low, which is partially explained by the fact that we calculate agreement over all languages together, rather than on the micro-level for individual languages. This is because there is no ground truth regarding the number of languages to annotate. Furthermore, a number of inconsistencies are due to ISO-639-3 variants (jap $\rightarrow$ jpn or ger $\rightarrow$ deu) and different codes for possibly ambiguous language mentions (‘Norwegian’ could be nor or nob). We iteratively resolve these issues to get our final list of languages per paper.
@@ 4 Results
4.1 Publication Overview
In total we retrieve 194 papers, of which 110 are found to contain a claim. The rest of this analysis is based on these 110 papers. Figure 2 shows the most common venues where papers with a claim have been published. The top six venues are all high-ranking NLP conferences, with SIGMORPHON being the first workshop at the seventh spot. The AI-centered venues only have a small number of papers with claims; ICLR and AAAI both have two, ICML one, and the rest none. In total, 38 papers introduce a dataset, most of which are published at ACL and LREC since 2020 (Figure 1 and 2). [^5] Details are listed in Appendix D, Table 4.
@@ 4.2 Language Overview
Four of the papers that claim typological diversity do not mention which languages they use at all. The other papers use a median of 11 languages ($Q1\!=\!8$, $Q3\!=\!18$), with a minimum of 2 and a maximum of 90. In total, the papers use 315 unique languages, of which 160 are used just once. Figure 3 shows the distribution of the number of languages used per paper.
English is the most used language (63 papers), followed by German (60), Russian (58) and Finnish (57). [^6] The long tail of this distribution is partially included in Appendix C, Figure 8. The paper that uses the most languages is Vylomova et al. (2020), which contains 90. Figure 4 shows where these languages are primarily spoken according to the World Atlas of Language Structures (WALS) Dryer and Haspelmath (2013). We observe a skew towards languages spoken in Europe, where multiple languages are used in more than 50 papers, while not a single language from the Americas is included in more than 10 papers.
@@ 4.3 Justifications
Still, most other papers justify their typological diversity in a ‘post-hoc’ way. They generally do not mention initial language selection considerations. Rather, they mention how diverse the sample is after it has been created. The XTREME-R dataset Ruder et al. (2021), which we will discuss further in §6, is exemplary of this. They mention ‘diversity indices’ that cover family, as well as typology and compare theirs to other datasets, but do not detail how or if these were used in sampling.
@@ 5 Typological Analysis
From Section 4.2), it follows that there is a geographical skew, where languages from certain areas are over-represented (Figure 4). This is not surprising, as it is often more feasible to select languages for which there are existing resources and typological descriptions, than languages for which these have to be gathered from scratch. This constitutes bibliographic bias (Rijkhoff and Bakker, 1998). However, we cannot draw conclusions about typological diversity on this basis, as typology is not equal to geography, nor phylogeny (Cysouw, 2013). Therefore, in this section, we analyze the typological diversity according to metrics that take into account syntactic features, and the absolute included typological feature values. Ponti et al. (2020) previously quantified language diversity with typology (URIEL), geography and family indices. These metrics do not take into account the distance between pairs of languages, and thus do not provide insights into the independence of languages in a sample. Additionally, no code or clear descriptions are provided, which limits their use.
Our metrics [^7] Formal definitions are provided in Appendix B. include pairwise language distances (§5.1) and absolute typological feature value inclusion (§5.2) across papers. Lastly, we look into the relationship between the number of included languages and typological diversity, to help guide future language selection efforts (§5.3).
@@ 5.1 Mean Language Distance
We first approximate the typological diversity of each paper’s language selection with lang2vec (Littell et al., 2017). This toolkit provides language distances that are calculated based on language vectors from URIEL, a resource that contains information from a range of typological databases, including WALS. We choose lang2vec specifically, because its aggregated information enables analysis across databases, and because its vectorized format mitigates issues stemming from incomplete feature coverage in typological databases.
For each paper in our survey that claims a typologically diverse language set and that is covered in lang2vec (312 / 315 unique languages), we calculate the mean pairwise syntactic distance (MPSD) for the included languages. That is, we take the average of all pre-computed syntactic lang2vec distances for each pair of languages in a given paper’s selection. We only take into account languages that have at least 5% coverage in the URIEL vectors. We do this to make sure the very low coverage languages do not distort the results. This affects 31 papers in our dataset, resulting in a total of 242 unique languages. The top-right plot in Figure 5 shows the distribution of MPSDs across papers. [^8] To illustrate this: the distance between Danish and Norwegian (same family, same genus) is 0.22, Danish-Spanish (same family, different genus) is 0.59 and Danish-Japanese (different family, different genus) is 0.69. We compare our typology-based approximations with mean geographical and genetic pairwise distances. Interestingly, the genetic distance is much higher and less spread out than the syntactic distance. This emphasizes that genealogical sampling does not by default ensure typological diversity.
The typological distances vary considerably, with outliers on either side (Figure 5). In papers claiming typological diversity, the minimum MPSD is found in Goel et al. (2022), who use English, French and Spanish. The maximum MPSD is found for Vania et al. (2019) their selection of North Sámi, Galician, and Kazah.